OpenStudy (anonymous):

Need Help will give MEDAL! Jesse can run around the track at his school in 64.3 seconds. The track is 370 meters long. Assuming Jesse could keep up the same pace, which number sentence could be used to find m, the number of meters Jesse could run in 321.5 seconds? A. m = (321.5 x 64.3) ÷ 370 B. m = 321.5 ÷ 64.3 C. m = (321.5 ÷ 64.3) x 370 D. m = (321.5 + 64.3) + 370

OpenStudy (anonymous):

@DullJackel09

OpenStudy (anonymous):

@linda3

OpenStudy (anonymous):

@Squirrels

OpenStudy (anonymous):

@baileypole

OpenStudy (anonymous):

Need help badly!!!!!!!!!!!! :(

OpenStudy (anonymous):

I think it is C.

OpenStudy (anonymous):

Do you guys have any idea on what the answer is?

OpenStudy (anonymous):

@linda3 do you think you know what the answer is? I think it is C.

OpenStudy (anonymous):

@hello1213 I need help BADLY!!!!!!!!!!!! :( HAAAAAAA Somebody please help!

OpenStudy (linda3):

i dont know i'm sorry :(

OpenStudy (anonymous):

OK do you know anyone who could help me?

OpenStudy (linda3):

@iambatman

OpenStudy (anonymous):

sorry no one is on who I am fans with... @terenzreignz is usually good but he is not on.

OpenStudy (anonymous):

@iambatman Please HELP ME!!!!!!!!!!!!! :(

OpenStudy (anonymous):

@texaschic101 Do you think you know the answer?

OpenStudy (anonymous):

I think the answer is C.

OpenStudy (dulljackel09):

Idk sorry

OpenStudy (dulljackel09):

Go with your answer then

OpenStudy (texaschic101):

Its A. I would have set it up a little different...I would have made a proportion....but basically answer A is doing the same thing that my proportion would do.

OpenStudy (anonymous):

K

OpenStudy (anonymous):

Thank you. :)

OpenStudy (texaschic101):

Let me show you... 370/64.3 = 321.5/x cross multiply 370x = (64.3)(321.5) x = (64.3)(321.5) / 370 your basically doing the same thing as answer A

OpenStudy (texaschic101):

sure thing :)